When a body is projected from the ground with an initial velocity'u'with a certain angle with respect to the ground, the magnitude of change in velocity from t=0 to t=0 secs is:

Grade:12th Pass

2 Answers

vikas askiitian expert
13 years ago

initial velocity = Vi = ucos@ (i) + usin@j

 

after t time ,

 (vertical comp of velocity ) Vy = (usin@-gt) (j)

 

 

horizontal component remains same ..

                   Vx = ucos@

 

final velocity (Vf) = (ucos@) (i) + (usin@-gt) (j)

 

change in velocity = Vi- Vf = dV = gt (j)

 

dV = gt  (change in velocity after time t)

approve if u like this ans

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Horizontal component of velocity remains same...         

So, Vx = ucos@

 

final velocity (Vf) = (ucos@) (i) + (usin@-gt) (j)

 

change in velocity = Vi- Vf = dV = gt (j)

 

dV = gt  (change in velocity after time t)

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi